OpenStudy (anonymous):

Can someone double-check this derivative for me? In comments, may take a minute to typeset.

OpenStudy (anonymous):

\[ \Large f(x)=\sin(x^{\sin(x^{\sin x})}) \]

OpenStudy (anonymous):

ick

OpenStudy (anonymous):

I'm getting the following result: \[\Large f'(x)=\cos(x^{\sin(x^{\sin x})})\cdot x^{\sin(x^{\sin x})}\\\cdot \Large \left[ \cos(x^{\sin x})\cdot x^{\sin x}\cdot \left(\cos x \log x + \frac{\sin x}{x}\right)\cdot \log x +\frac{\sin(x^{\sin x})}{x} \right] \] The answer key in the book has something slightly different, not sure if it's a typo or if I'm confused.

OpenStudy (anonymous):

\[\frac{d}{dx}x^{\sin(x)}=x^{\sin(x)}(\cos(x)\ln(x)+\frac{\sin(x)}{x})\] and then the chain rule

OpenStudy (anonymous):

Pretty sure it's a typo, it has \((\cos x \log x + \dfrac{\cos x}{x})\), which doesn't make any sense.

OpenStudy (anonymous):

no that is not right unless i messed up gotta run, good luck

OpenStudy (anonymous):

My first step was rewriting it as follows: \[\huge f(x)=\sin(e^{\sin (e^{\sin x \log x})\log x}) \]

OpenStudy (anonymous):

I'm almost positive the answer in the book is wrong, but I just want to double-check before moving on.

OpenStudy (asnaseer):

you can use wolframalpha to check your steps: http://www.wolframalpha.com/input/?i=d%2Fdx+of+sin%28x^%28sin%28x^%28sin%28x%29%29%29%29%29